OpenStudy (anonymous):

Use the inverse matrix method to solve the following systems of equations. 2x+y=5 2x-y=-1

OpenStudy (anonymous):

just add and get \(4x=4\) making \(x=1\)

OpenStudy (anonymous):

and y = 3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can u help me with another one ?

OpenStudy (anonymous):

depends on if i know it or not

OpenStudy (anonymous):

go ahead and post

OpenStudy (anonymous):

x+y+z=7 2x-6z=2 3x-2y+5z=13

OpenStudy (anonymous):

oh three ick i would cheat, this is donkey work or maybe someone else can show you how to do it step by step

OpenStudy (anonymous):

lol . wow !

OpenStudy (anonymous):

if you just need the answer, do this http://www.wolframalpha.com/input/?i=x%2By%2Bz%3D7+%2C2x-6z%3D2%2C+3x-2y%2B5z%3D13

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

otherwise you can eliminate z from the first and third, then get two equations with two unknowns it is totally doable, but takes a while and is a pain